OpenStudy (anonymous):

the total number of democrats and republicans in the us house of representivess during the 108th session was 434. there were 24 more republicans than democrats. how many members of session of each party were there? the total number of democrats and republicans in the us house of representivess during the 108th session was 434. there were 24 more republicans than democrats. how many members of session of each party were there? @Mathematics

OpenStudy (anonymous):

step by step plz

OpenStudy (anonymous):

(434/2) - (24/2) = 205 of one, 205+24 = 229 of the other.

OpenStudy (hoblos):

let # democrats=x & # republicans=y => x+y =434 and we have y=x+24 by solving the system we get => x= 205 y=229

OpenStudy (anonymous):

i dont get it

OpenStudy (hoblos):

which part u didn't get?